OpenStudy (anonymous):

8. Write these products as sums: a. cos(2x)sin(5x) b. sin(3x)cos(2x) c. cos(7x)cos(2x)

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

http://www.thattutorguy.com/trigonometry/power-reducing-formulas-for-sine-cosine/ For example: \[\sin (5 x) \cos (2 x)=\frac{1}{2} (\sin (3 x)+\sin (7 x)) \]